If I would have purchased 11 articles for Rs. 10 and sold all at the rate 10 for Rs. 11. The profit percent would have been;

A. 10
B. 11
C. 21
D. 100

Kindly, tell me what does the question mean and how to solve the question, please.

Thnk you.
1
Expert's answer
2020-08-13T17:18:24-0400

cost price of 11 articles = Rs. 10

cost price of 1 article = Rs. "\frac{10}{11}"

selling price of 10 articles = Rs . 11

selling price of 1 article = Rs. "\frac{11}{10}"

since selling price > cost price , there is profit

profit per article = selling price - cost price

= Rs. "\frac{11}{10}" - Rs."\frac{10}{11}"


= Rs."\frac{121-100}{110}"


= Rs."\frac{21}{110}"

Profit on 11 articles = "\frac{21}{110}"x11 = Rs. 2.1

profit % = "\frac{profit}{costprice}"x 100%

= "\frac{2.1}{10}"x100%


=21%

correct option is C .